OpenStudy (anonymous):

What are the solutions of the system? y = x^2 + 4.6x + 4.7 y = 3.3x - 3.7 A. (–3.47, 3.48) and (2.17, –15.17) B. (–3.47, –15.17) and (–2.17, –10.88) C. (–3.47, –15.17) and (2.17, 3.48) D. no solution

OpenStudy (zzr0ck3r):

you can solve or plug each value in and check

OpenStudy (anonymous):

If we set them equal (y=y) We get this: x^2 + 4.6x + 4.7= 3.3x - 3.7 Now we can substract 3.3x both sides x^2+1.3x+4.7=-3.7 Now we add 3.7 both sides: x^2+1.3x+8.4=0 There we can use the quadratic formula. Remember it?